ROTECTED VIEW Be cavefulles from the Internet can contain viuses. Unless you need to ndit, its safler to stay in Prodected View.Enable Editing A clothing retailer gives you 10 points for every $1 spent, and you can trade in 2,350 points for a $15 gft certificate. Eddie Bauer Shopping Spree Iteres Cost Total Cost Points L Polo Shirt a) The spreadsheet lists purchases and their prices. Create formulas which compute the cost of each item and the points you get from each item. 5 135.00 5219.00 19.99 s 75.00 4 Merrell Shoes Snowline Jacket 7 Cloud Layer Fleece b.) Compute the total cost of the order and the total number of points from the order Full Dip Hoodie 0 Wrinkle Free Oxfor $ 70.00 55.00 Totals: c-) Compute how many $15 gft certificates you will get from this order and the total value of these gift certificates. Note: you cannot get a fractional gift certificate so use the INT function which gives the integer part of a number (effectively it rounds down). Format with er decimal places Gift Certificates: Value of Gift Certificates: Money SpenValue of Gift cestificates d) What is the ratio of money spent to the value of the gift certificates? Enter a formula in cell 121 for this ratio with the second quantity scaled to one. 82 EddieB
Answer #1

a) Formula

Total Cost for polo shirt - =c4*d4

Total points - =E4*10

b) Calculation of total cost and total points

Items Quantity Cost Total cost Total points
Polo shirt 3 $50 $150 $1500
Dress shirt 4 $75 $300 $3000
Khakis 2 $59.99 $119.98 $1199.8
Merell shirt 1 $135 $135 $1350
Snowline Jacket 1 $219 $219 $2190
Belt 1 $19.99 $19.99 $199.9
Cloud layer fleece 2 $75 $150 $1500
Sweater 3 $80 $240 $2400
Full zip hoodie 1 $70 $70 $700
Wrinkle Free oxford 5 $55 $275 $2750
Total $1678.97 $16789.7

c) Calculation of gift certificates and its value

One gift certificate of $15 for 2550 points

Formula for number of gift certificates - =INT(F15/2550)

Gift certificate = 6 Certificate

Formula for Value of gift certificates - =F17*$15

Value of gift certificates = 6*$15 = $90

d) Ratio of money spent to value of gift certificates

Formula for ratio = E15:F18

Ratio = $1678.97:$90